solve for x please
3x - 5(3x - 7) = 2(x + 9) + 45

Answer by tutorcecilia(2144)

3x - 5(3x - 7) = 2(x + 9) + 45
3x-15x+35=2x+18+45
-12x+35=2x+63
-12x+35-2x-63=0
-14x-28=0
-14x=28
-14x/-14=28/-14
x=-2
.
.Check by pluggining (x=-2) back into the original equation]